OpenStudy (anonymous):

find the distance between the given line and point 1.) x = 7y - 3 (1, -1) Long method 2.) x-y = 1 (6,5) Short method 3.) 3x -4y -10 = 0 (-1 , 3) any method help.. ;D

OpenStudy (anonymous):

hi try graphing the number 1) 7y=x+3...y=(x+3)/7

OpenStudy (anonymous):

do you know the distance formula?

OpenStudy (anonymous):

using the point(1,-1), at x=1, y=(x+3)/7=4/7 this is point(1,4/7) distance formula d=sqrt[(x2-x1)^2+(y2-y1)^2] d=sqrt[(1-1)^2+(-1-4/7)^2] d=-11/7

OpenStudy (anonymous):

2.) x-y = 1 (6,5) set x=6 6-y=1 6-1=y 5=y, therefore this is point(6,5) distance formula d=sqrt[(x2-x1)^2+(y2-y1)^2] distance formula d=sqrt[(6-6)^2+(5-5)^2] d=0 3.) 3x -4y -10 = 0 (-1 , 3) ,,set x=-1 3(-1)-4y-10=0 -3-10=-13=4y -13/4=y use point(-1,-13/4) and (-1,3) d=sqrt[(-1+1)^+(-13/4 -3)^2 d=-25/4

OpenStudy (anonymous):

absolute value nang AX BY +C all over 2a (short method) square root of (X1+X2)(Y1+Y2) (long method)